# Physics-Informed Neural Systems for the Simulation of EUV Electromagnetic Wave Diffraction from a Lithography Mask

### 16 Mar 2026 v1

@@ 3.2 Waveguide Method
$\psi_{mn}=\exp\left(-i\kappa_{x}mx-i\kappa_{y}ny\right)$
From the system of Equations (2) we have the following equations
@@
$-\sum\limits_{m,n}(\kappa_{x}^{2}m^{2}+\kappa_{y}^{2}n^{2})B_{mn}\psi_{mn}$ $+k_{0}^{2}\sum\limits_{l,p}\sum\limits_{m,n}\varepsilon_{lp}B_{mn}$
$-i\sum\limits_{l,p}\sum\limits_{m,n}\hat{\varepsilon}^{(y)}_{lp}\left(\kappa_{x}mC_{mn}-\kappa_{y}nB_{mn}\right)\psi_{(m+l)(n+p)}=k_{z}^{2}\sum\limits_{m,n}B_{mn}\psi_{mn}\psi_{(m+l)(n+p)},$
@@ 3.3 Physics-Informed Neural Networks
$\left.\int\limits_{-L_{x}/2}^{L_{x}/2}\int\limits_{-L_{y}/2}^{L_{y}/2}\left(\frac{\partial{\bf u}}{\partial z}+i\tilde{k}_{z;mn}{\bf u}\right)\frac{1}{\psi_{mn}}dxdy\right|_{z=z_{\max}}=0,$
$\left.\int\limits_{-L_{x}/2}^{L_{x}/2}\int\limits_{-L_{y}/2}^{L_{y}/2}\left(\frac{\partial{\bf u}}{\partial z}-i\tilde{k}_{z;mn}{\bf u}\right)\frac{1}{\psi_{mn}}dxdy\right|_{z=z_{\min}}=0,$ (23)
where the first two conditions are the periodicity condition of the solution, the last two are the Sommerfeld conditions at the upper ($z_{\max}$) and lower ($z_{\min}$) boundaries of the domain under consideration. $\left\{{\bf x}_{i}^{(bc)}\right\}^{N_{bc}}_{i=1}$ and $\left\{{\bf x}_{i}^{(r)}\right\}^{N_{r}}_{i=1}$ are sets of points corresponding to boundary condition domain and PDE domain, respectively. These points can be the vertices of a fixed mesh or can be randomly sampled at each iteration of a gradient descent algorithm. All required gradients w.r.t. input variables ($\bf x$) and network parameters $\theta$ can be efficiently computed via automatic differentiation [10] with algorithmic accuracy, which is defined by the accuracy of computation system.
@@ 4.1 Validation on Test Problems
To evaluate the accuracy of the solutions provided by numerical solvers (FreeFem++, WG) and solvers based on artificial neural networks, we have tested the methods on three 2D problems with known analytical solutions: (1) plane wave propagation in a homogeneous medium, (2) reflection from a single interface (Figure 6a), and (3) reflection from a dielectric layer (Figure 6b). As the main performance metrics for the considered solvers, we evaluated the relative ${\mathbb{L}}_{2}$ error (see [7] and appendix A) and the time to provide a solution (the time to calculate a solution for numerical solver or the inference time for neural networks). In addition, training times are marked separately for neural networks.
@@
For the wavelength of 11.2 nm, the top three layers of the mask follow the same configuration as described above: an absorber consisting of a TaBO layer (10 nm) and a TaBN layer (10 nm for PINN, 60 nm for WGNO). Since a standardized solution for mirrors at this wavelength has not yet been established and related research is still ongoing [4, 21, 23], we considered several potential mirror designs for further experiments. Figure 10 shows the reflectivity as a function of the angle of incidence for Mo/Be, Ru/Be, and Ru/Be/Sr mirrors, each consisting of 60 layers with thicknesses specified in Table 3. As can be seen, the Ru/Be/Sr mirror exhibits the highest reflectivity. Consequently, this structure was chosen as the mask substrate (see Figure 2(c)). For the PINN-based solver, a single layer of this structure was utilized, whereas for the WGNO, the mirror consisted of a 31-layer Ru/Be/Sr. The dielectric permittivities of TaBO and TaBN are $\varepsilon=0.906{-}i0.064$ and $\varepsilon=0.909{-}i0.060$ respectively. As in the previous mask case, the absorbers had a hole with width $L_{x}/2$ so the permittivities of the absorbers are described by Eqn. 29. The angle of incidence of the wave is $6^{\circ}$. Figures 10 and 12 show the comparisons of the calculated $E_{y}$ field, normalized to the amplitude of the incident wave, predicted by PINN and WGNO, respectively, for the wavelength 11.2 nm.
@@
Figures 13 and 14 show cross-sections of the field calculated by our WGNO compared to the reference solution at 13.5 nm and 11.2 nm. The agreement is excellent, demonstrating the capability of our method to handle the full 3D problem with high fidelity.
@@ Appendix A Relative ${\mathbb{L}}_{2}$ error
Figures 1518 present a comparison between the exact solutions of the test problems and those obtained using FreeMem++, the waveguide method, PINN, and WGNO. The selected test cases include: plane wave propagation in a homogeneous medium, reflection from a single interface (Figure 6a), and reflection from a dielectric layer (Figure 6b).

### 17 Mar 2026 v2

@@ 3.2 Waveguide Method
$\psi_{mn}=\exp\left(-i\kappa_{x}mx-i\kappa_{y}ny\right).$
From the system of Equations (3.2) we have the following equations
@@
$-\sum\limits_{m,n}(\kappa_{x}^{2}m^{2}+\kappa_{y}^{2}n^{2})B_{mn}\psi_{mn}$ $+k_{0}^{2}\sum\limits_{l,p}\sum\limits_{m,n}\varepsilon_{lp}B_{mn}\psi_{(m+l)(n+p)}$
$-i\sum\limits_{l,p}\sum\limits_{m,n}\hat{\varepsilon}^{(y)}_{lp}\left(\kappa_{x}mC_{mn}-\kappa_{y}nB_{mn}\right)\psi_{(m+l)(n+p)}=k_{z}^{2}\sum\limits_{m,n}B_{mn}\psi_{mn},$
@@ 3.3 Physics-Informed Neural Networks
$\left.\int\limits_{-L_{x}/2}^{L_{x}/2}\int\limits_{-L_{y}/2}^{L_{y}/2}\left(\frac{\partial{\bf u}}{\partial z}+i{k}_{z;mn}{\bf u}\right)\frac{1}{\psi_{mn}}dxdy\right|_{z=z_{\max}}=0,$
$\left.\int\limits_{-L_{x}/2}^{L_{x}/2}\int\limits_{-L_{y}/2}^{L_{y}/2}\left(\frac{\partial{\bf u}}{\partial z}-i{k}_{z;mn}{\bf u}\right)\frac{1}{\psi_{mn}}dxdy\right|_{z=z_{\min}}=0,$ (23)
where the first two conditions are the periodicity condition of the solution, the last two are the Sommerfeld conditions at the upper ($z_{\max}$) and lower ($z_{\min}$) boundaries of the domain under consideration. $\left\{{\bf x}_{i}^{(bc)}\right\}^{N_{bc}}_{i=1}$ and $\left\{{\bf x}_{i}^{(r)}\right\}^{N_{r}}_{i=1}$ are sets of points corresponding to boundary condition domain and PDE domain, respectively. These points can be the vertices of a fixed mesh or can be randomly sampled at each iteration of a gradient descent algorithm. All required gradients w.r.t. input variables ($\bf x$) and network parameters ($\bm{\theta}$}) can be efficiently computed via automatic differentiation [10] with algorithmic accuracy, which is defined by the accuracy of computation system.
@@ 4.1 Validation on Test Problems
To evaluate the accuracy of the solutions provided by numerical solvers (FreeFem++, WG) and solvers based on artificial neural networks, we have tested the methods on three 2D problems with known analytical solutions: (1) plane wave propagation in a homogeneous medium, (2) reflection from a single interface (Figure 6 (a)), and (3) reflection from a dielectric layer (Figure 6 (b)). As the main performance metrics for the considered solvers, we evaluated the relative ${\mathbb{L}}_{2}$ error (see [7] and appendix A) and the time to provide a solution (the time to calculate a solution for numerical solver or the inference time for neural networks). In addition, training times are marked separately for neural networks.
@@
For the wavelength of 11.2 nm, the top two layers of the mask follow the same configuration as described above: an absorber consisting of a TaBO layer (10 nm) and a TaBN layer (10 nm for PINN, 60 nm for WGNO). Since a standardized solution for mirrors at this wavelength has not yet been established and related research is still ongoing [4, 21, 23], we considered several potential mirror designs for further experiments. Figure 10 shows the reflectivity as a function of the angle of incidence for Mo/Be, Ru/Be, and Ru/Be/Sr mirrors, each consisting of 60 layers with thicknesses specified in Table 3. As can be seen, the Ru/Be/Sr mirror exhibits the highest reflectivity. Consequently, this structure was chosen as the mask substrate (see Figure 2 (c)). For the PINN-based solver, a single layer of this structure was utilized, whereas for the WGNO, the mirror consisted of a 31-layer Ru/Be/Sr. The dielectric permittivities of TaBO and TaBN are $\varepsilon=0.906{-}i0.064$ and $\varepsilon=0.909{-}i0.060$ respectively. As in the previous mask case, the absorbers had a hole with width $L_{x}/2$ so the permittivities of the absorbers are described by Eqn. (29). The angle of incidence of the wave is $6^{\circ}$. Figures 11 and 12 show the comparisons of the calculated $E_{y}$ field, normalized to the amplitude of the incident wave, predicted by PINN and WGNO, respectively, for the wavelength 11.2 nm.
@@
Figures 13 and 14 show cross-sections of the field calculated by our WGNO compared to the reference solution at 13.5 nm and 11.2 nm, respectively. The agreement is excellent, demonstrating the capability of our method to handle the full 3D problem with high fidelity.
@@ Appendix A Relative ${\mathbb{L}}_{2}$ error
Figures 1518 present a comparison between the exact solutions of the test problems and those obtained using FreeMem++, the waveguide method, PINN, and WGNO. The selected test cases include: plane wave propagation in a homogeneous medium, reflection from a single interface (Figure 6 (a)), and reflection from a dielectric layer (Figure 6 (b)).